1.     Suppose that under the United States-Canada Air Quality Agreement, both countries agree to a combined 30 percent SO2 abatement standard. Further assume that the following SO2 abatement cost functions have been estimated by each country:

 

                                              TACUS = 500 + 1.5(AUS)2

                                              MACUS = 3 AUS

                                              TACCAN = 1000 + 3(ACAN)2

                                              MACCAN = 6 ACAN

A.     If the countries implement a uniform abatement standard, find the resulting values of TAC and MAC for each nation. Based on these values, is there an economic incentive the two nations to participate in an emissions trading program? Explain. (10 marks)

B.     Assuming a trading program is enacted, find the cost savings associated with a cost-effective abatement solution? (12 marks)

C.     What must be the price of a tradeable permit to achieve the cost-effective abatement allocation? (3 marks)



1
Expert's answer
2021-12-08T19:18:24-0500

(A)

If uniform abatement gets implemented then each country abates half of the total targeted abatement standard which is 30% / 2 = 15% each.

 

At such standard the TAC and MAC of each country US:

TACus = 500 + 1.5(A)2

      = 500 + 1.5(15)2

      = 837.5

 

MACus = 3(A)

       = 3(15) 

       = 45

 

Calculate the TÁC and MAC of Country CAN:

TACcan = 1000 + 3(A)2

       = 1000 + 3(15)2

       = 1675

 

MACcan = 6(A)

        = 6(15)

        = 90

 

Hence at uniform abatement standard the total abatement cost and marginal abatement cost of US are $837.5 millions and $45 millions respectively whereas the total abatement cost and marginal abatement cost of Canada are $1675 millions and millions respectively.

 

As it is seen that the marginal cost of both countries are not equal at uniform abatement standard it implies that one country incurs high cost than other in same abatement thus inefficiency occurs and produces incentives for countries to participate into emission trading program.

Step 2

(B)

If a trading program gets enacted then the result of such program will be equalization of marginal abatement cost of both countries.

MACus  = MACcan

3Aus   =   6Acan

Aus = 2Acan  .......................(1)

 

And combined abatement standard is 30%

that is,

Acan + Aus = 30

30 - Acan = Aus ......................(2)

 

From equation (1) and (2)

30 - Acan = 2 Acan

Acan = 10%

 

And thus,

Aus = 2(10) = 20%

At these levels of abatement total costs are as follows:


Earlier total combined cost of abatement was 837.5+1675 = $2512.5 millions and now it is 1100 +1300 = $2400 millions hence the total cost that gets saved equals $2512.5- $2400 $112.5 millions.

Step 3

(C)

At cost effective level Country C abates five percent less and U.S abates five percent more. Considering a permit equal to abate one percent the price of each permit should not be more than $6 as the marginal cost of abatement of Country C is $6 . On the other hand, it should be at least $3 as the marginal cost of abatement of U.S. is $3.